OpenStudy (anonymous):

Can somebody Please give me the steps to do: x^11y^3 --------- x^11y

OpenStudy (fibonaccichick666):

So, when you divide like bases with exponents, do you add, subtract, multiply, or divide the exponents?

OpenStudy (anonymous):

I've no clue.

OpenStudy (fibonaccichick666):

Ok, what if you had \(x^4x^6\), what would that simplify to?

OpenStudy (anonymous):

I think it would be x^10

OpenStudy (fibonaccichick666):

That would be correct. What did you do?

OpenStudy (anonymous):

I kept the base the same since X is the same as X and just added the exponents.

OpenStudy (fibonaccichick666):

That is very good now, if you have \(\frac{x^6}{x^4}\), what would you do?

OpenStudy (anonymous):

Im sorry, but I can't read what that says.

OpenStudy (fibonaccichick666):

\[\frac{x^6}{x^4}\]

OpenStudy (anonymous):

I would do xxxxxx ------ xxxx

OpenStudy (anonymous):

and then cross out the x's that match so that I am left with two x's

OpenStudy (fibonaccichick666):

COntinue

OpenStudy (anonymous):

so then it would be x to the second?

OpenStudy (phi):

*** I would do xxxxxx ------ xxxx **** That is beautiful thinking. The last step is to notice the short cut. 6 x's up top 4 x's in the bottom. 6-4 = 2

OpenStudy (fibonaccichick666):

Ok, now take a look at your original exponents, what would you do to get \[6 ? 4=2\]

OpenStudy (anonymous):

6-4=2

OpenStudy (fibonaccichick666):

So in general, if you have division of like bases you _______. If you have multiplication you _______.

OpenStudy (anonymous):

If its division of like bases you keep them the same and if you have multiplication I dont know.

OpenStudy (fibonaccichick666):

You just told me, that sentence is the same as \(x^4x^6\) and \[ {\frac{x^6}{x^4}}\]

OpenStudy (anonymous):

oh I get it.

OpenStudy (fibonaccichick666):

What did you do to the exponents

OpenStudy (anonymous):

you just add the exponents and keep the bases the same if there already the same.

OpenStudy (fibonaccichick666):

Technically that is correct for both, but For which were you implying? (might fall under semantics)

OpenStudy (anonymous):

ok.

OpenStudy (fibonaccichick666):

So which one do you add?

OpenStudy (anonymous):

the exponents.

OpenStudy (fibonaccichick666):

but for which function? (multiplication or division)

OpenStudy (anonymous):

both

OpenStudy (fibonaccichick666):

uhm, well depends on if you've learned something yet if that is correct

OpenStudy (fibonaccichick666):

have you seen \[x^{-2}\]

OpenStudy (fibonaccichick666):

can you re write that another way if you have?

OpenStudy (anonymous):

yes, I think If I rewrite x^-2 It would be -x times -x

OpenStudy (fibonaccichick666):

Ok, no that is not correct. So yes, normally when you have exponents, they are the short hand way of multiplying that many of the same number, but when you have a negative exponent it means '1 over" so ie.

OpenStudy (fibonaccichick666):

\[x^n=x_1x_2....x_n\]\[x^{-n}=\frac{1}{x_1}\frac{1}{x_2}...\frac{1}{x_n}\]

OpenStudy (fibonaccichick666):

Do you understand that^?

OpenStudy (anonymous):

Not at all.

OpenStudy (fibonaccichick666):

oh k uhm let me try and think of an easier way.... uh personally, I can't, so try and read this http://www.purplemath.com/modules/exponent.htm

OpenStudy (fibonaccichick666):

this might be easier http://www.mathsisfun.com/algebra/exponent-laws.html, but both are good

OpenStudy (anonymous):

cannot find the file it says

OpenStudy (fibonaccichick666):

which?

OpenStudy (anonymous):

the mathisfun one.

OpenStudy (fibonaccichick666):

oh delete the comma at the end

OpenStudy (fibonaccichick666):

http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

ok Thank you for all of the help. you are very smart.

OpenStudy (fibonaccichick666):

You're trying to learn, that is all I can ask for.

OpenStudy (fibonaccichick666):

Thank you, so once you read through that, we can go over this problem. Just come up with some ideas on how to simplify it, ok?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

ok so my original problem was: x^11y^3 ------- x^11y so putting that into scientific notation would be xxxxxxxxxxx yyy --------------- xxxxxxxxxxx y wich would cancel out all the x's and leave it with just y^2 Is this correct?

OpenStudy (fibonaccichick666):

That is, but can you do it without writing it all out?

OpenStudy (anonymous):

No. I wasn't taught that there was another way do tell

OpenStudy (fibonaccichick666):

Ok that second link shows it. They are called the basic laws/rules

OpenStudy (fibonaccichick666):

so let m and n be any number

OpenStudy (anonymous):

oh,cool.

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

if you have \[x^mx^n=x^{m+n}\] Do you follow?

OpenStudy (anonymous):

I do.

OpenStudy (fibonaccichick666):

ok now here is where that fraction \(\frac{1}{x^n}=x^{-n}\] comes into play

OpenStudy (fibonaccichick666):

oops \[\frac{1}{x^n}=x^{-n}\]

OpenStudy (fibonaccichick666):

If you have \[ x^m x^{-n}= \frac {x^m} {x^{n}}= x^{m-n}\]

OpenStudy (fibonaccichick666):

ok, do you follow that

OpenStudy (anonymous):

yes.

OpenStudy (fibonaccichick666):

so now, you essentially have \[\frac{x^{11}}{x^{11}} \times \frac{y^3}{y}\]

OpenStudy (fibonaccichick666):

Using the rules I just taught you, can you simplify this?

OpenStudy (anonymous):

nope, I've got no clue. sorry eli5

OpenStudy (fibonaccichick666):

...ok if I tell you that for x, m=11 AND n=11 can you figure it out then?

OpenStudy (anonymous):

yup I think so, m multiplied by n and 11+11 ?

OpenStudy (fibonaccichick666):

no, look at the eq please.

OpenStudy (fibonaccichick666):

We have division

OpenStudy (anonymous):

whats the eq?

OpenStudy (anonymous):

I just don't understand it. Is it ok if I just write it out?

OpenStudy (fibonaccichick666):

\[\frac{x^{11}}{x^{11}} \times \frac{y^3}{y}\]

OpenStudy (fibonaccichick666):

Not really, you need to learn this. It will save you astronomical amounts of time

OpenStudy (anonymous):

alright.

OpenStudy (anonymous):

I shall learn.

OpenStudy (fibonaccichick666):

ie \[x^{100}x^{200}\]

OpenStudy (fibonaccichick666):

can you simplify that in less than 10 seconds?

OpenStudy (anonymous):

x^300

OpenStudy (fibonaccichick666):

ok, how did you do that?

OpenStudy (anonymous):

I kept the bases the same since the both are x and then I added the exponents together.

OpenStudy (fibonaccichick666):

good! now \[\frac{x^{200}}{x^{100}}\]

OpenStudy (anonymous):

200-100=100, so x^100?

OpenStudy (fibonaccichick666):

YES! now

OpenStudy (anonymous):

yay

OpenStudy (fibonaccichick666):

Now, \[\frac{x^{11}}{x^{11}}\]

OpenStudy (anonymous):

11-11=0 so x^0= just 1

OpenStudy (fibonaccichick666):

YES! now \[\frac{y^3}{y}\]

OpenStudy (anonymous):

3-0 ? if so then 3-0=3 so y^3

OpenStudy (fibonaccichick666):

why is it 0?

OpenStudy (fibonaccichick666):

As you just noted, anything to the zero power is 1

OpenStudy (fibonaccichick666):

but anything to the _____ power is itself

OpenStudy (anonymous):

oh so it's 3-1=2 so y^2

OpenStudy (fibonaccichick666):

YUP!

OpenStudy (fibonaccichick666):

Now, your eq. \[\frac{x^{11}}{x^{11}} \times \frac{y^3}{y}\]

OpenStudy (fibonaccichick666):

Can you simplify this using this method?

OpenStudy (anonymous):

11-11=0=1 3-1=2 so 1 times 2?

OpenStudy (fibonaccichick666):

you forgot to keep your bases

OpenStudy (anonymous):

oh yeah, woops. X*Y^2 I think that would be it?

OpenStudy (fibonaccichick666):

Stop you forgot something again